Physics Question: Please provide an answer with shown work!

Problem 9.12 Part A Friends throw snowballs at the wall of a 3.0 m × 6.0 m barn. The snowballs have mass 0.10 kg and hit the wall moving at an average speed of 6.0 m/s. They do not rebound. Determine the average pressure exerted by the snowballs on the wall if 60 snowballs hit the wall each second Express your answer to two significant figures and include the appropriate units. P Value Units Submit My Answers Give Up Incorrect; Try Again; 4 attempts remaining Provide Feedback Continue
Answer #1

A = 3 x 6 = 18 m^2


Impulse = change in momentum = 0.10 (6 - 0) = 0.60 kg m/s

force on wall = impulse imparted per second

= 0.60 x 60

= 36 kg m/ s^2


P = F / A = 36 / 18 = 2 N / m^2 Or Pa